# XiYan-SQL: A Multi-Generator Ensemble Framework for Text-to-SQL

### v1

@@ Abstract
To tackle the challenges of large language model performance in natural language to SQL tasks, we introduce XiYan-SQL, an innovative framework that employs a multi-generator ensemble strategy to improve candidate generation. We introduce M-Schema, a semi-structured schema representation method designed to enhance the understanding of database structures. To enhance the quality and diversity of generated candidate SQL queries, XiYan-SQL integrates the significant potential of in-context learning (ICL) with the precise control of supervised fine-tuning. On one hand, we propose a series of training strategies to fine-tune models to generate high-quality candidates with diverse preferences. On the other hand, we implement the ICL approach with an example selection method based on named entity recognition to prevent overemphasis on entities. The refiner optimizes each candidate by correcting logical or syntactical errors. To address the challenge of identifying the best candidate, we fine-tune a selection model to distinguish nuances of candidate SQL queries. The experimental results on multiple dialect datasets demonstrate the robustness of XiYan-SQL in addressing challenges across different scenarios. Overall, our proposed XiYan-SQL achieves the state-of-the-art execution accuracy of 89.65% on the Spider test set, 69.86% on SQL-Eval, 41.20% on NL2GQL, and a competitive score of 72.23% on the Bird development benchmark. The proposed framework not only enhances the quality and diversity of SQL queries but also outperforms previous methods.
@@ 1 Introduction
We present comprehensive evaluations on both relational and non-relational databases, specifically focusing on prominent systems such as SQLite, PostgreSQL, and nGQL. XiYan-SQL demonstrates remarkable performance across a range of benchmarks, achieving the state-of-the-art performance on the Spider [32], SQL-Eval, and NL2GQL [33] datasets with 89.65%, 69.86%, and 41.20% execution accuracy, respectively. In the context of the more challenging Bird [10] benchmark, XiYan-SQL also achieves a competitive score of 72.23%. The impressive results achieved on various challenging NL2SQL benchmarks not only validate the effectiveness of our approach but also demonstrate its significant potential for broader applications in NL2SQL translation tasks. XiYan-SQL can be accessed from https://bailian.console.aliyun.com/xiyan. We also release the source code for connecting to the database and building M-Schema at https://github.com/XGenerationLab/M-Schema.
@@ 7.2 Bird Results
We compare the performance of different NL2SQL methods on Bird development benchmark in Table 3. XiYan-SQL achieves a competitive accuracy of 72.23%, much higher than 57.95% of GPT-4o. CHASE-SQL [17] framework employs multiple chain-of-thought prompting techniques to generate candidates, and subsequently implements a binary voting mechanism among 21 candidates, achieving an accuracy of 73.14%. XiYan-SQL yields a competitive performance by voting among only 5 candidates.
We also observe that a significant number of the leading methods on the bird learderborad are based on prompt engineering techniques. It suggests the immense potential of large-scale models and the importance of carefully designed prompts in optimizing model performance. The SFT based method, ExSL + Granite-34B-Code, secures the second position with an accuracy of 72.43%. This notable performance indicates that, smaller-sized models are indeed capable of generating complex SQL queries effectively through advanced training techniques. XiYan-SQL integrates the methodologies of SFT and ICL to balance the test time and the overall performance of the system.
Table 2: Performance comparison of different NL2SQL methods on Bird development benchmark.
Method EX(%)
CHASE-SQL + Gemini [17] 73.14
ExSL + granite-34b-code 72.43
Insights AI 72.16
AskData + GPT-4o 72.03
OpenSearch-SQL, v2 + GPT-4o 68.90
CHESS [25] 68.31
PURPLE + RED + GPT-4o 68.12
Distillery + GPT-4o [16] 67.21
MCS-SQL [25] 63.36
SuperSQL [8] 58.50
SFT CodeS-15B [9] 58.47
GPT-4o 57.95
TA-SQL + GPT-4 [21] 56.19
DAIL-SQL [5] 54.76
XiYan-SQL 72.23
@@ 7.6.1 M-Schema
We conduct ablation study on the Bird development benchmark to present the impact of different schema representations on end-to-end SQL generation performance. To demonstrate the generalization ability of our proposed M-Schema, we use fout powerful LLMs as NL2SQL generators, DeepSeek [14], Claude 3.5 Sonnet  [^3] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 1.5 Pro and GPT-4o [1]. As shown in Table 6, all four models have performance improvements using M-Schema as the representation of database schema compared to DDL Schema, with an average increase of 2.03%. Although M-schema is similar to MAC-SQL Schema in structure, GPT-4o and Claude 3.5 Sonnet show 0.65% and 0.78% improvements, respectively. While DeepSeek and Gemini 1.5 have slight accuracy decreases of 0.13% and 0.26%. The experimental results indicate that M-Schema is a better representation than DDL Schema and MAC-SQL Schema and demonstrates powerful generalizability.
@@ 7.6.2 Schema Linking
XiYan-SQL All (five candidates) 72.23 -
@@ 7.6.3 Candidate Generation and Selection
To evaluate the effectiveness and impact of candidate generation and selection, we conduct various ablation studies on XiYan-SQL. Table 8 presents the performance of XiYan-SQL when certain components are dropped, highlighting their significance in achieving high-quality performance. The "XiYan-SQL All" method achieves an accuracy of 71.51% by utilizing three candidates, of which two are generated from two distinct fine-tuned SQL generators (as described in Section 5.1), while one is produced by the ICL SQL generator with GPT-4o (as presented in Section 5.2). For the candidate generator, there is a significant decrease in the performance of XiYan-SQL when the fine-tuned candidate generators are removed, further indicating that our generator is capable of generating high-quality and diverse candidate SQL queries. Similarly, the removal of the ICL generator and Refiner also leads to a decline in performance. Additionally, concerning candidate selection, we observe that when the selection model is not employed, relying solely on self-consistency for candidate selection, XiYan-SQL’s performance decreases by approximately three percentage points. This finding underscores the effectiveness of our proposed method. Finally, when the number of SQL candidates is increased to five, the accuracy of XiYan-SQL can further reach 72.23%.

### v2

@@ Abstract
To tackle the challenges of large language model performance in natural language to SQL tasks, we introduce XiYan-SQL, an innovative framework that employs a multi-generator ensemble strategy to improve candidate generation. We introduce M-Schema, a semi-structured schema representation method designed to enhance the understanding of database structures. To enhance the quality and diversity of generated candidate SQL queries, XiYan-SQL integrates the significant potential of in-context learning (ICL) with the precise control of supervised fine-tuning. On one hand, we propose a series of training strategies to fine-tune models to generate high-quality candidates with diverse preferences. On the other hand, we implement the ICL approach with an example selection method based on named entity recognition to prevent overemphasis on entities. The refiner optimizes each candidate by correcting logical or syntactical errors. To address the challenge of identifying the best candidate, we fine-tune a selection model to distinguish nuances of candidate SQL queries. The experimental results on multiple dialect datasets demonstrate the robustness of XiYan-SQL in addressing challenges across different scenarios. Overall, our proposed XiYan-SQL achieves the state-of-the-art execution accuracy of 75.63% on Bird benchmark, 89.65% on the Spider test set, 69.86% on SQL-Eval, 41.20% on NL2GQL. The proposed framework not only enhances the quality and diversity of SQL queries but also outperforms previous methods.
@@ 1 Introduction
We present comprehensive evaluations on both relational and non-relational databases, specifically focusing on prominent systems such as SQLite, PostgreSQL, and nGQL. XiYan-SQL demonstrates remarkable performance across a range of benchmarks, achieving the state-of-the-art performance on the Spider [32], SQL-Eval, and NL2GQL [33] datasets with 89.65%, 69.86%, and 41.20% execution accuracy, respectively. In the context of the more challenging Bird [10] benchmark, XiYan-SQL also reaches the top accuracy of 75.63%. The impressive results achieved on various challenging NL2SQL benchmarks not only validate the effectiveness of our approach but also demonstrate its significant potential for broader applications in NL2SQL translation tasks. XiYan-SQL can be accessed from https://bailian.console.aliyun.com/xiyan. We also release the source code for connecting to the database and building M-Schema at https://github.com/XGenerationLab/M-Schema.
@@ 7.2 Bird Results
We compare the performance of different NL2SQL methods on Bird benchmark in Table 3. XiYan-SQL reaches the top of Bird leaderboard with an accuracy of 75.63%, outperforming the second place of 0.84%. CHASE-SQL [17] framework employs multiple chain-of-thought prompting techniques to generate candidates, and subsequently implements a binary voting mechanism among 21 candidates, achieving an accuracy of 74.70%. XiYan-SQL yields a competitive performance by voting among only 5 candidates.
We also observe that a significant number of the leading methods on the bird learderborad are based on prompt engineering techniques. It suggests the immense potential of large-scale models and the importance of carefully designed prompts in optimizing model performance. The SFT based method, ExSL + Granite-34B-Code, secures the fourth position with an accuracy of 73.17%. This notable performance indicates that, smaller-sized models are indeed capable of generating complex SQL queries effectively through advanced training techniques. XiYan-SQL integrates the methodologies of SFT and ICL to balance the test time and the overall performance of the system.
Table 2: Performance comparison of different NL2SQL methods on Bird benchmark.
Method EX(Dev) EX(Test)
CHASE-SQL + Gemini [17] 74.46 74.79
DSAIR + GPT-4o 74.32 74.12
ExSL + granite-34b-code 72.43 73.17
AskData + GPT-4o 72.03 72.39
OpenSearch-SQL, v2 + GPT-4o 69.30 72.28
Distillery + GPT-4o [16] 67.21 71.83
CHESS [25] 68.31 71.10
Insights AI 72.16 70.26
PURPLE + RED + GPT-4o 68.12 70.21
MCS-SQL [25] 63.36 65.45
SuperSQL [8] 58.50 62.66
SFT CodeS-15B [9] 58.47 60.37
GPT-4o 57.95 -
TA-SQL + GPT-4 [21] 56.19 59.14
DAIL-SQL [5] 54.76 57.41
XiYan-SQL 73.34 75.63
@@ 7.6.1 M-Schema
We conduct ablation study on the Bird development benchmark to present the impact of different schema representations on end-to-end SQL generation performance. To demonstrate the generalization ability of our proposed M-Schema, we use four powerful LLMs as NL2SQL generators, DeepSeek [14], Claude 3.5 Sonnet  [^3] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 1.5 Pro and GPT-4o [1]. As shown in Table 6, all four models have performance improvements using M-Schema as the representation of database schema compared to DDL Schema, with an average increase of 2.03%. Although M-schema is similar to MAC-SQL Schema in structure, GPT-4o and Claude 3.5 Sonnet show 0.65% and 0.78% improvements, respectively. While DeepSeek and Gemini 1.5 have slight accuracy decreases of 0.13% and 0.26%. The experimental results indicate that M-Schema is a better representation than DDL Schema and MAC-SQL Schema and demonstrates powerful generalizability.
@@ 7.6.2 Schema Linking
XiYan-SQL All (five candidates) 73.34 -
@@ 7.6.3 Candidate Generation and Selection
To evaluate the effectiveness and impact of candidate generation and selection, we conduct various ablation studies on XiYan-SQL. Table 8 presents the performance of XiYan-SQL when certain components are dropped, highlighting their significance in achieving high-quality performance. The "XiYan-SQL All" method achieves an accuracy of 71.51% by utilizing three candidates, of which two are generated from two distinct fine-tuned SQL generators (as described in Section 5.1), while one is produced by the ICL SQL generator with GPT-4o (as presented in Section 5.2). For the candidate generator, there is a significant decrease in the performance of XiYan-SQL when the fine-tuned candidate generators are removed, further indicating that our generator is capable of generating high-quality and diverse candidate SQL queries. Similarly, the removal of the ICL generator and Refiner also leads to a decline in performance. Additionally, concerning candidate selection, we observe that when the selection model is not employed, relying solely on self-consistency for candidate selection, XiYan-SQL’s performance decreases by approximately three percentage points. This finding underscores the effectiveness of our proposed method. Finally, when the number of SQL candidates is increased to five, the accuracy of XiYan-SQL can further reach 73.34%.